# Metric Differential Privacy at the User-Level

### v1

@@ 1.1.1. Mechanism Design
First, we study how to release linear queries $V\tilde{K}_{G}$, where $\tilde{K}_{G}\in\mathbb{R}^{\mathcal{X}}$ is the normalized representation of the global dataset $K_{G}$ and $V\in\mathbb{R}^{d\times\mathcal{X}}$ is a real-valued matrix with bounded entries. While computing the sensitivity of a linear query is easy under user-level DP, proving a sensitivity under $d_{\textsf{EM}}$-DP is quite challenging. Specifically, it requires analysis of a coupling between two possible datasets, along with a stronger assumption that $V$ is “Lipschitz” in a sense, rather than just being bounded. To this end, we first prove the following bound:
@@
(Informal version of Thm. 4.1): The sensitivity of $V\tilde{K}_{G}$ is upper bounded by
$\max_{K,K^{\prime}}\frac{\|V\tilde{K}_{G}-V\tilde{K}_{G}^{\prime}\|}{d_{\textsf{EM}}(\tilde{K}_{G},{\tilde{K}_{G}}^{\prime})}\leq\max_{x,x^{\prime}\in\mathcal{X}}\frac{\|V[x]-V[x^{\prime}]\|}{d_{\mathcal{X}}(x,x^{\prime})},$
where the notation $V[x]$ indicates the column of $V$ indexed by $x$.
Using the above result, we show that the sensitivity of $V$, which is a maximum over the space of all datasets, can be reduced to a Lipschitz property of $V$ that is much easier to compute. In Sec.  6.1, we show that a special class of linear queries, which we call linear embedding queries, satisfies the above mentioned Lipschitzness and can provide provably better utility than user-level DP.
@@
We design a mechanism for performing itemwise queries on the entire dataset $K$. Our approach is to simply apply a private mechanism $\mathcal{A}$ to each item $k_{i}\in K$ and release the set of noisy outputs $\{\mathcal{A}(k_{i})\}$ after shuffling them. Here $\mathcal{A}$ can be an arbitrary mechanism satisfying $(\alpha,0)$-$d_{\mathcal{X}}$ DP which makes our mechanism completely general-purpose (see Sec. 4.2 for some concrete examples of $\mathcal{A}$). Here we consider the bounded data setting since the size of $K$ is revealed. The main technical novelty lies in providing a tight privacy analysis of the above mechanism. Specifically, prior work shows that the above mechanism satisfies bounded $(m\alpha,0)$-$d_{\textsf{EM}}$-DP (Fernandes et al., 2019) by using the interplay between couplings and privacy via composition. However, we show that composition is not the right tool for tight privacy analysis since it does not take into that the output of our mechanism is an unordered list, i.e., the $\mathcal{A}(k_{i})$s are released in a random arbitrary order. Instead, we generalize a tight result from privacy amplification by shuffling (Feldman et al., 2022) to metric DP.
@@
(Informal version of Thm. 4.3) Suppose that $\mathcal{A}:\mathcal{X}\rightarrow\mathcal{Y}$ is an $\alpha$-$d_{\mathcal{X}}$ DP algorithm with respect to $d_{\mathcal{X}}$. Let $(x_{1},\ldots,x_{m})\in\mathcal{X}^{m}$ be a dataset. Then, releasing $\mathsf{Shuffle}(\mathcal{A}(x_{1}),\ldots,\mathcal{A}(x_{m}))$ satisfies $(O(\alpha\sqrt{me^{\alpha}\ln(m/\delta)}),\delta e^{\alpha})$-$d_{\textsf{EM}}$ DP.
@@
1.1.2. Extending $d_{\textsf{EM}}$-DP to the Unbounded Setting
We start our mechanism designs by considering the bounded data settings in both the local and central models of privacy (See Table 1) as this enables easier privacy analysis (Sec. 4). However, the bounded setting might be restrictive in practice as it cannot support usecases where users have different amounts of data, or the data sizes are also private. To this end, we extend $d_{\textsf{EM}}$-DP to the more general unbounded setting. We show that when user data is relatively homogeneous (such as, when it is i.i.d.), the privacy analysis of the unbounded setting may be reduced to the bounded setting.
Specifically, in Sec.  5 we create a black-box projection mechanism which projects any unbounded dataset onto a dataset where each user contributes a fixed, predefined amount of data. This enables running any bounded $d_{\textsf{EM}}$-DP mechanism on the projected data. Our projection mechanism samples a fixed number of dataset items with replacement from each user. The privacy analysis follows by showing that the $d_{\textsf{EM}}$ between any two datasets remains relatively unchanged by sampling, up to a small additive factor as determined by the Chernoff’s bound.
One caveat is that the introduced additive factor necessitates a slight adjustment to the privacy semantics of $d_{\textsf{EM}}$-DP. Instead of protecting any change of $d_{\textsf{EM}}$ distance $d$ with a privacy parameter $d\alpha$, we consider a small threshold $r$ such that all changes less than $r$ are protected with a uniform parameter $r\alpha$. In essence, this privacy guarantee provides $d_{\textsf{EM}}$-DP at the granularity of units of $d_{\textsf{EM}}$ distance $r$. We refer to this notion as discrete user-level $d_{\textsf{EM}}$-DP (Def. 5.1) and have the following result:
@@ 1.1.2. Extending $d_{\textsf{EM}}$-DP to the Unbounded Setting
(Informal version of Thm. 5.3) Suppose that for $n$ users, $\mathcal{M}$ is a mechanism which satisfies $(\alpha,\delta)$-bounded $d_{\textsf{EM}}$-DP. The algorithm which, given arbitrary user datasets $K_{1},\ldots,K_{n}$, takes $s$ i.i.d. samples from each $K_{i}$ and then applies $\mathcal{M}$ on each of the sampled data items, satisfies $(\alpha r,\delta,r)$-discrete $d_{\textsf{EM}}$-DP for all $r\geq\frac{2\ln(1/\delta)}{s}$.
@@
1.1.3. Demonstrating Improvements Over User-level DP
We compare the privacy and utility of our proposed $d_{\textsf{EM}}$-DP mechanisms with baseline mechanisms satisfying user-level DP. Specifically, in Sec.  6.1, we study a special type of linear query called linear embedding queries and in Sec.  6.2, we study problem of private frequency estimation. For simplicity, we consider the bounded data setting.
@@ 1.1.3. Demonstrating Improvements Over User-level DP
  • $\alpha=\varepsilon$: Since [^3] If $\alpha\leq\varepsilon$, then user-level DP is strictly weaker than $d_{\textsf{EM}}$-DP; the more appropriate baseline is to use $\alpha=\varepsilon$. we assume $d_{\mathcal{X}}$ is normalized, we always have $d_{\textsf{EM}}\leq 1$. Thus, in this case $(\varepsilon,\delta)$-$d_{\textsf{EM}}$-DP implies $(\varepsilon,\delta)$-user-level DP. However, any pair of input $K,K^{\prime}$ such that $d_{\textsf{EM}}(\tilde{K},\tilde{K}^{\prime})<1$ the privacy protection of $d_{\textsf{EM}}$-DP is actually stronger.
@@
Now, we interpret the theoretical error bounds for linear embedding queries. From Table 2(a), the error for releasing a $d$-dimensional linear embedding query under user-level DP is $O(\frac{d}{\varepsilon n})$, while it is $O(\frac{d}{\alpha n})$ for $d_{\textsf{EM}}$-DP. When $\alpha=\varepsilon$, these utilities are identical, but $d_{\textsf{EM}}$ DP offers stronger privacy. When $\alpha>\varepsilon$, then the utility of $d_{\textsf{EM}}$-DP is higher than that of user-level DP, with the the two guarantees offering differing privacy semantics. Thus, in both cases, there is a clear benefit of using $d_{\textsf{EM}}$-DP. These observations are the same in the local model.
Finally, for frequency estimation in the local model, Table 2(b) shows that the error of user-level DP is $O(\sqrt{\frac{k^{2}\ln(m/\delta)}{n\varepsilon^{2}}})$, while it is $O(\sqrt{\frac{k^{3}}{n\alpha^{2}}\max\{\ln(\frac{m}{\delta},\alpha\}}$ for $d_{\textsf{EM}}$ DP. For constant $\varepsilon$ and $\alpha\geq\varepsilon^{2}\frac{k}{\ln(m/\delta)}$, the utility is improved. In the central model, the error of the user-level DP algorithm is $O(\frac{k}{n\varepsilon})$ while it is $O(\frac{k^{3/2}}{n\alpha}\sqrt{\max\{\ln(\frac{m}{\delta}),\alpha\}})$ for $d_{\textsf{EM}}$-DP. The algorithm under $d_{\textsf{EM}}$-DP has the added benefit that it can be implemented in the shuffle model of privacy, which requires less trust and parallels prior work in the shuffle model (Feldman et al., 2022). There is a utility improvement for $\alpha\geq\varepsilon^{2}k$. When $\varepsilon\leq\alpha\leq\varepsilon^{2}k$, we leave it as an interesting open problem whether $d_{\textsf{EM}}$-DP can offer utility improvements over user-level DP.
@@
Laplace Mechanism $(\varepsilon,0)$-user level DP Central, Bounded $O(\frac{d}{\varepsilon n})$ (Lemma 6.3) $d_{\textsf{EM}}-DP$ gives same utility but stronger privacy $\alpha=\varepsilon$; $d_{\textsf{EM}}-DP$ gives better utility but different privacy for $\alpha>\varepsilon$.
@@
PrivEMDItemWise $(\alpha,\delta)$-$d_{\textsf{EM}}$-DP Central, Bounded  [^4] This algorithm works in the shuffle model, which requires less trust than the central model. $O\left(\frac{k^{3/2}}{n\alpha}\sqrt{\max\left\{\ln(\frac{m}{\delta}),\alpha\right\}}\right)$ (Corollary 6.8)
(b) Comparison of $d_{\textsf{EM}}$-DP to user-level DP for frequency estimation in the setting defined in Sec.  6.2. $k$ is the domain size $|\mathcal{X}|$.
Table 2. Summary of theoretical utility guarantees, assuming there are $n$ users who hold datasets of size $m$.
2. Background
2.1. Differential Privacy
Intuitively, DP is a property of a mechanism which ensures that its output distribution remains insensitive to changes in the data of a single individual. The standard DP guarantee, which is also know as item-level DP, considers each user $U_{i}$ to contribute only a single item $x_{i}\in\mathcal{X}$ to a global dataset, i.e., $K_{i}=x_{i}$. In this paper, we consider differential privacy at the user level. We start by considering the local model:
Definition 2.1 (Unbounded User-level Local DP (Acharya et al., 2023)).
@@ 2.1. Differential Privacy
(1) $\Pr[\mathcal{M}(K)=O]\leq e^{\varepsilon}\Pr[\mathcal{M}(K^{\prime})=O]+\delta.$
@@
Definition 2.2 (Unbounded User-level Central DP (Liu et al., 2023)).
@@
(2) $\Pr[\mathcal{M}(K_{G})=O]\leq e^{\varepsilon}\Pr[\mathcal{M}(K_{G}^{\prime})=O]+\delta.$
Note that there is no restriction in the sizes of the datasets $\{K_{i}\},i\in[n]$ in the above definition.
Next, we define metric DP ($d_{\mathcal{X}}$-DP) that enables the privacy guarantee to depend on the $d_{\mathcal{X}}$ distance between the pair of inputs. We start by introducing it at the item-level (so we consider changing an item $x\in\mathcal{X}$ to another item $x^{\prime}\in\mathcal{X}$). For simplicity, we consider the local model, so the mechanism acts on just a single item:
Definition 2.3 (Local $d_{\mathcal{X}}$-DP (Alvim et al., 2018)).
@@
$\Pr[\mathcal{M}(x)=O]\leq e^{\alpha d_{\mathcal{X}}(x,x^{\prime})}\Pr[\mathcal{M}(x^{\prime})=O].$
We replace the traditional privacy parameter $\varepsilon$ with $\alpha$ in the above definition, because $\varepsilon$ in Definitions 2.1 and 2.2 is a unitless parameter while $\alpha$ has the inverse unit of $d_{\mathcal{X}}$.
2.2. Earth-Mover’s Distance
Notations. We denote the set of all possible datasets as $\mathcal{X}^{*}$. We will also view a dataset $K\in\mathcal{X}^{*}$ as a probability distribution defined by its normalized histogram $\tilde{K}$. To do so, let $\Delta^{\mathcal{X}}\subseteq\mathbb{R}^{\mathcal{X}}$ denote the probability simplex indexed by $\mathcal{X}$—i.e. the set of all vectors $\langle v_{x}\rangle_{x\in\mathcal{X}}$ such that $v_{x}\geq 0$ and $\sum_{x\in\mathcal{X}}v_{x}=1$. For a dataset $K$, $\tilde{K}\in\Delta^{\mathcal{X}}$ then denotes the probability distribution defined by $K$, meaning $\tilde{K}[x]=\frac{\text{Num. occurrences of $x$ in $K$}}{|K|}$. A natural way to extend the notion of distance from items in $\mathcal{X}$ to distributions in $\Delta^{\mathcal{X}}$ is to use the Earth-Mover’s (or $1$-Wasserstein) distance (Givens and Shortt, 1984), which we now define. For a joint distribution $C(x_{1},x_{2})\in\Delta^{\mathcal{X}\times\mathcal{X}}$, let $C_{x_{1}}(x_{2})$ denote the distribution conditioned on observing $x_{1}$, and let $C_{1}(x_{1})$ denote the marginal distribution of $x_{1}$. We define $C_{x_{2}}(x_{1})$ and $C_{2}(x_{2})$ similarly.

### v2

@@ 1.1.1 Mechanism Design
First, we study how to release linear queries $F\tilde{K}$, where $F\in\mathbb{R}^{d\times|\mathcal{X}|}$ is a real-valued matrix with bounded entries. While computing the global sensitivity of such a query is easy under user-level DP, proving a sensitivity bound under $d_{\textsf{EM}}$-DP is considerably more complex. It involves making a stronger Lipschitz assumption about the points in $F$, and then leveraging this property to transport $\tilde{K}$ onto $\tilde{K}^{\prime}$ to compute an upper bound on how much $F\tilde{K}$ can change by. To this end, we first prove the following bound:
@@
(Informal version of Theorem 4.1): The sensitivity of $F\tilde{K}$ is upper bounded by
$\max_{K,K^{\prime}}\frac{\|F\tilde{K}-F\tilde{K}^{\prime}\|}{d_{\textsf{EM}}(\tilde{K},{\tilde{K}}^{\prime})}\leq\max_{x,x^{\prime}\in\mathcal{X}}\frac{\|F[x]-F[x^{\prime}]\|}{d_{\mathcal{X}}(x,x^{\prime})},$
where the notation $F[x]$ indicates the column of $F$ indexed by $x$.
Using the above result, we show that the sensitivity of $F$, which is a maximum over the space of all datasets, can be reduced to the Lipschitzness of $F$, which is simpler to bound. Sensitivity analysis for standard DP typically only requires $F$ to be bounded. However, in the case of $d_{\textsf{EM}}$-DP, we need a stronger assumption – $F$ must also be Lipschitz. In Section 4.1, we demonstrate that several commonly used queries, such as average, region, and similarity queries, are indeed Lipschitz.
@@
We design a mechanism for performing itemwise queries on the entire dataset $K$. For now we consider bounded DP, so $|K|$ is known in advance. Our approach is simple – apply a private mechanism $\mathcal{A}$ to each item $x_{j}\in K$ and then release the set of noisy outputs $\{\mathcal{A}(x_{j})\}$ after shuffling them. Here $\mathcal{A}$ can be an arbitrary mechanism satisfying $(\alpha,0)$-$d_{\mathcal{X}}$ DP which makes our mechanism completely general-purpose (see Section 4.2 for some concrete examples of $\mathcal{A}$). The main technical novelty lies in the privacy analysis of the above mechanism. While one can use composition to show this release satisfies $O(\alpha m)$-$d_{\textsf{EM}}$ DP Fernandes et al. (2019), this is far from being tight. As discussed in Section 4.2, composition is not the right tool for tight privacy analysis since it does not account for the fact that the output of our mechanism is an unordered list, i.e., the $\mathcal{A}(x_{j})$s are released in a random order. Instead, we leverage privacy amplification by shuffling (Feldman et al., 2022) to prove a general shuffling result and adapt it for $d_{\textsf{EM}}$-DP.
@@
(Informal version of Theorem 4.3) Suppose that $\mathcal{A}:\mathcal{X}\rightarrow\mathcal{Y}$ is an $\alpha$-$d_{\mathcal{X}}$ DP algorithm with respect to $d_{\mathcal{X}}$. Let $(x_{1},\ldots,x_{m})\in\mathcal{X}^{m}$ be a dataset. Then, releasing $\mathsf{Shuffle}(\mathcal{A}(x_{1}),\ldots,\mathcal{A}(x_{m}))$ satisfies $(O(\alpha\sqrt{me^{\alpha}\ln(m/\delta)}),\delta e^{\alpha})$-$d_{\textsf{EM}}$ DP.
@@
1.1.2 Extending $d_{\textsf{EM}}$-DP to the Unbounded Setting
We start our mechanism designs by considering the bounded data setting in both the local and central models of privacy (see Table 1) as this enables easier privacy analysis (Section 4). However, the bounded setting might be restrictive in practice as it cannot support usecases where users have different amounts of data, or the data sizes are also private. To this end, we extend $d_{\textsf{EM}}$-DP to the more general unbounded setting. We achieve this through a general reduction that converts any bounded $d_{\textsf{EM}}$-DP mechanism into an unbounded one while treating the mechanism as a black box. If the data from each user is relatively homogeneous, such as being i.i.d., then the utility of the mechanism will be preserved.
Our reduction uses a projection mechanism that projects any dataset onto one with a fixed size, without significantly increasing the $d_{\textsf{EM}}$ distance. The projection we use is sampling with replacement. Intuitively, this is a smooth projection because we can view sampling from two datasets $K,K^{\prime}$ in terms of a coupling between them, and show that the sampled points can also be coupled with expected cost given by $d_{\textsf{EM}}(K,K^{\prime})$. Using Bernstein’s inequality, we show convergence to $d_{\textsf{EM}}(K,K^{\prime})$, up to a small additive factor.
One caveat is that the introduced additive factor necessitates a slight adjustment to the privacy semantics of $d_{\textsf{EM}}$-DP. Whereas $d_{\textsf{EM}}$-DP protects a change of $d_{\textsf{EM}}$-distance $d$ with an effective privacy parameter $\alpha d$, for all $d>0$, our modified definition includes $r$ as a fixed parameter, and states that all changes of $d_{\textsf{EM}}\leq r$ are protected uniformly with parameter $\alpha r$. We may view this as a discretization of the $d_{\textsf{EM}}$ by rounding it up to the nearest multiple of $r$. We refer to this notion as $(\alpha r,\delta,r)$-discrete user-level $d_{\textsf{EM}}$-DP (Def. 5.1). This privacy guarantee is weaker than $d_{\textsf{EM}}$-DP only for changes of $d_{\textsf{EM}}<r$—for bigger changes, the two definitions are equivalent up to factors of $2$ using group privacy. Typically, $r$ is small and the practical difference between the definitions is negligible. Our reduction satisfies the following:
@@ 1.1.2 Extending $d_{\textsf{EM}}$-DP to the Unbounded Setting
(Informal version of Theorem 5.3) Suppose that for $n$ users, $\mathcal{M}$ is a mechanism which satisfies $(\alpha,\delta)$-bounded $d_{\textsf{EM}}$-DP. The algorithm which, given arbitrary user datasets $K_{1},\ldots,K_{n}$, takes $s$ i.i.d. samples from each $K_{i}$ and then applies $\mathcal{M}$ on each of the sampled data items, satisfies $(\alpha r,\delta,r)$-discrete $d_{\textsf{EM}}$-DP (in the central model) for all $r\geq\frac{2\ln(1/\delta)}{s}$.
@@
1.1.3 Demonstrating Improvements Over User-level DP
Finally, we evaluate the benefit of using the more nuanced privacy semantics of $d_{\textsf{EM}}$-DP over standard user-level DP by comparing the privacy and utility of our proposed mechanisms with baselines. Specifically, in Sec. 6.1, we study a special type of linear query called linear embedding queries and in Sec. 6.2, we study private frequency estimation. For simplicity, we consider the bounded data setting.
@@ 1.1.3 Demonstrating Improvements Over User-level DP
  • $\alpha=\varepsilon$: Since we assume $d_{\mathcal{X}}$ is normalized, we always have $d_{\textsf{EM}}\leq 1$. Thus, in this case $(\varepsilon,\delta)$-$d_{\textsf{EM}}$-DP implies $(\varepsilon,\delta)$-user-level DP. However, any pair of input $K,K^{\prime}$ such that $d_{\textsf{EM}}(\tilde{K},\tilde{K}^{\prime})<1$ the privacy protection of $d_{\textsf{EM}}$-DP is actually stronger. (Note that if $\alpha\leq\varepsilon$, then user-level DP is strictly weaker than $d_{\textsf{EM}}$-DP; the more appropriate baseline is to use $\alpha=\varepsilon$.)
@@
Now, we interpret the theoretical error bounds for linear embedding queries. From Table 2(a), the error for releasing a $d$-dimensional linear embedding query under user-level DP is $O(\frac{d}{\varepsilon n})$, while it is $O(\frac{d}{\alpha n})$ for $d_{\textsf{EM}}$-DP. When $\alpha=\varepsilon$, these utilities are identical, but $d_{\textsf{EM}}$-DP offers stronger privacy. When $\alpha>\varepsilon$, then the utility of $d_{\textsf{EM}}$-DP is higher than that of user-level DP, with the the two guarantees offering differing privacy semantics. Thus, in both cases, there is a clear benefit of using $d_{\textsf{EM}}$-DP. These observations are the same in the local model.
Finally, for frequency estimation in the local model, Table 2(b) shows that the error of user-level DP is $O(\sqrt{\frac{k^{2}\ln(m/\delta)}{n\varepsilon^{2}}})$, while it is $O(\sqrt{\frac{k^{3}}{n\alpha^{2}}\max\{\ln(\frac{m}{\delta}),\alpha\}}$ for $d_{\textsf{EM}}$-DP. For constant $\varepsilon$ and $\alpha\geq\varepsilon^{2}\frac{k}{\ln(m/\delta)}$, the utility is improved. In the central model, the error of the user-level DP algorithm is $O(\frac{k}{n\varepsilon})$ while it is $O(\frac{k^{3/2}}{n\alpha}\sqrt{\max\{\ln(\frac{m}{\delta}),\alpha\}})$ for $d_{\textsf{EM}}$-DP. The algorithm under $d_{\textsf{EM}}$-DP has the added benefit that it can be implemented in the shuffle model of privacy, which requires less trust and parallels prior work in the shuffle model Feldman et al. (2022). There is a utility improvement for $\alpha\geq\varepsilon^{2}k$. When $\varepsilon\leq\alpha\leq\varepsilon^{2}k$, we leave it as an interesting open problem whether $d_{\textsf{EM}}$-DP can offer utility improvements over user-level DP.
@@
$K$-norm Mechanism $(\varepsilon,0)$-user level DP Central, Bounded $O(\frac{d}{\varepsilon n})$ (Lemma 6.3) $d_{\textsf{EM}}$-DP gives same utility but stronger privacy for $\alpha=\varepsilon$; $d_{\textsf{EM}}$-DP gives better utility but different privacy for $\alpha>\varepsilon$.
@@
For random linear queries, this algorithm also has best-known error among all $(\varepsilon,\delta)$-user-level DP algorithms.
@@
PrivEMDItemWise $(\alpha,\delta)$-$d_{\textsf{EM}}$-DP Central, Bounded $O\left(\frac{k^{3/2}}{n\alpha}\sqrt{\max\left\{\ln(\frac{m}{\delta}),\alpha\right\}}\right)$ (Corollary 6.8)
This algorithm works in the shuffle model, which requires less trust than the central model.
This algorithm also has best-known error among all $(\varepsilon,\delta)$-user-level DP algorithms.
(b) Comparison of $d_{\textsf{EM}}$-DP to user-level DP for frequency estimation in the setting defined in Section  6.2. $k$ is the domain size $|\mathcal{X}|$.
Table 2: Summary of theoretical utility guarantees, assuming there are $n$ users who hold datasets of size $m$.
2 Background
2.1 Differential Privacy
Intuitively, DP is a property of a mechanism which ensures that its output distribution remains insensitive to changes in the data of a single individual. The standard DP guarantee, which is also know as item-level DP, considers each user $U_{i}$ to contribute only a single item $x_{i}\in\mathcal{X}$ to a global dataset, i.e., $K_{i}=x_{i}$. Instead we consider user-level DP, where each $K_{i}$ is itself a multiset of elements from $\mathcal{X}$. We denote the set of all multisets of elements from $\mathcal{X}$ as $\mathcal{X}^{*}$. In the local model, our privacy definition is then:
Definition 2.1 (Unbounded User-level Local DP Acharya et al. (2023)).
@@ 2.1 Differential Privacy
$\Pr[\mathcal{M}(K)=O]\leq e^{\varepsilon}\Pr[\mathcal{M}(K^{\prime})=O]+\delta.$ (1)
@@
Definition 2.2 (Unbounded User-level Central DP Liu et al. (2023)).
@@
$\Pr[\mathcal{M}(K_{G})=O]\leq e^{\varepsilon}\Pr[\mathcal{M}(K_{G}^{\prime})=O]+\delta.$ (2)
Note that there is no restriction on the sizes of the datasets $\{K_{i}\},i\in[n]$ in the above definition.
Next, we define metric DP that enables the privacy guarantee to depend on a metric $d_{\mathcal{X}}$ between the pair of inputs. We start by introducing it at the item-level (i.e., we consider changing only one item $x\in\mathcal{X}$ to another item $x^{\prime}\in\mathcal{X}$). For simplicity, we consider the local model, so the mechanism acts on just a single item:
Definition 2.3 (Local $d_{\mathcal{X}}$-DP Alvim et al. (2018)).
@@
$\Pr[\mathcal{M}(x)=O]\leq e^{\alpha d_{\mathcal{X}}(x,x^{\prime})}\Pr[\mathcal{M}(x^{\prime})=O]+\delta.$
We replace the traditional privacy parameter $\varepsilon$ with $\alpha$ in the above definition, because $\varepsilon$ in Definitions. 2.1 and 2.2 is a unitless parameter while $\alpha$ has the inverse unit of $d_{\mathcal{X}}$.
2.2 Earth-Mover’s Distance
Notations. We view datasets as multisets of elements from $\mathcal{X}$. We will also view a dataset $K\in\mathcal{X}^{*}$ as a probability distribution defined by its normalized histogram $\tilde{K}$. To do so, let $\Delta^{\mathcal{X}}\subseteq\mathbb{R}^{\mathcal{X}}$ denote the probability simplex indexed by $\mathcal{X}$—i.e. the set of all vectors $\langle v_{x}\rangle_{x\in\mathcal{X}}$ such that $v_{x}\geq 0$ and $\sum_{x\in\mathcal{X}}v_{x}=1$. For a dataset $K$, $\tilde{K}\in\Delta^{\mathcal{X}}$ then denotes the probability distribution defined by $K$, meaning $\tilde{K}[x]=\frac{\text{Num. occurrences of $x$ in $K$}}{|K|}$. The earth-mover’s (or $1$-Wasserstein) distance Givens and Shortt (1984) is defined as follows. For a joint distribution $C(x_{1},x_{2})\in\Delta^{\mathcal{X}\times\mathcal{X}}$, let $C_{x_{1}}(x_{2})$ denote the distribution conditioned on observing $x_{1}$, and let $C_{1}(x_{1})$ denote the marginal distribution of $x_{1}$. We define $C_{x_{2}}(x_{1})$ and $C_{2}(x_{2})$ similarly.